Okay, the following question has me stumped, Geometry: If the roof of a home is 7m from peak to eave and the attic space is 3m high at the peak, how long does each of the pieces of insulation need to be? (Insulation is 40cm wide rolls that are cut to fit between the rafters in the attic) Round the answer to the nearest tenth.
Answer by checkley75(3666)

WHAT YOU HAVE IS A RIGHT TRIANGLE WITH THE RIGHT SIDE=3M & THE HYPOTENUS IS 7M.
NOW WWE NEED TO FIND THE OTHER SIDE (BASE) USING THE PYTHAGOREAN THERUM A^2+B^2=C^2 WHERE A & B ARE THE 2 SIDES & C IS THE HYPOTENUSE.
3^2+B^2=7^2
9+X^2=49
X^2=49-9
X^2=40
X=SQRT40
X=6.3M IS THE BASE MEASUREMENT.